Hi, I'm having trouble figuring out this question:
Determine the domain of the variable "x" in each expression.
1.) 4 / x-5
2.) x / x+4

Answer by jim_thompson5910(35256)

I'll do problem 1 to get you started.

I'm assuming the entire denominator is x-5. If so, then set the denominator equal to 0 and solve for x

x-5 = 0
x-5+5 = 0+5
x+0 = 5
x = 5

So if x = 5, then the denominator x-5 is equal to 0

Recall that you cannot divide by zero. So we must exclude x = 5 from the denominator to avoid division by zero errors.

Therefore, the domain is the set of all real numbers BUT x cannot equal 5. You can pick any other number you want for x.